OpenStudy (anonymous):

integral of 1/(x^2+x)

OpenStudy (anonymous):

\[\int\limits_{}^{} 1/ (x^2+x) \]

OpenStudy (lgbasallote):

how about factoring first \[\int \frac{1}{x(x+1)}\] i believe you can use partial fractions here yes?

OpenStudy (anonymous):

i understand but is it possible to use intyegration by parts as well?

OpenStudy (lgbasallote):

hmm do you mean \[\int \frac 1x \times \frac{1}{x+1}\] interesting...

OpenStudy (lgbasallote):

well i can assure you it's gonna be complicated IF possible

OpenStudy (lgbasallote):

let u = 1/x du = -1/x^2 dv = 1/x+1 v = ln (x+1) do you see it? the complication i mean

OpenStudy (lgbasallote):

you're gonna have \[\frac{\ln (x+1)}{x} + \int \frac{\ln(x+1)}{x^2}dx\] how do you suppose you can integrate that

OpenStudy (anonymous):

okat thank you i see your point